Solve the following equations:
(5/2) x = 25/4...........

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

First we have to multiply both the side by 2,
= (5x/2) × 2 = (25/4) × 2
= 5x = (25/2)
Now,
We have to divide both the side by 5,
Then we get,
= 5x/5 = (25/2)/5
= x = (25/2) × (1/5)
= x = (5/2)

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
We can solve this question as follows :
(5/2)x = 25/4
5x / 2 = 25/4
By cross multiplication,
20x = 50
X = 50/20
X = 5/2
 
Hope it helps!
Thank you ✌️